### 12 Mar 2026 v1

@@ 5.1 Baseline methods
We select COLMAP [33] as the baseline to represent the classical optimization-based structure-from-motion method. Since COLMAP only generates sparse point models, we limit our comparison to camera pose evaluation when using COLMAP. In experiments, we set the known camera intrinsics for COLMAP.
We also evaluate against recent state-of-the-art feed-forward models: Fast3R [43], VGGT [37], and a memory-efficient version of VGGT FastVGGT [34]. These methods demonstrate strong performance in scenarios where traditional methods often fail. Finally, we include CUT3R [39] as a baseline due to its specific design for video processing, while it also supports unstructured image inputs. When running CUT3R, we sequentially feed frames from each camera video (i.e., camera-1 followed by camera-2) to fully leverage its temporal processing capabilities.
@@ 5.2 Metrics
We evaluate depth quality using Absolute Relative Depth (Abs Rel) and Delta accuracy [24, 46]($\delta<1.25$, i.e., the percentage of predicted depths within a 1.25-factor of true depth). Per-frame depth evaluation is conducted by scaling the predicted depth and offset to the ground truth.
@@ 5.3 Datasets
In addition to our self-collected MultiCamRobolab dataset, we also evaluate our method against the MultiCamVideo-Dataset [2]. MultiCamVideo-Dataset is a multi-camera synchronized video dataset rendered in simulation using Unreal Engine, which includes synchronized multi-camera videos and their corresponding camera trajectories. Each video clip has 81 frames with 10 different camera perspectives. In our evaluation, we randomly extract 50 clips and 3 random cameras for each clip in the MultiCamVideo-Dataset. All images are cropped to $512\times 384$, and trajectories are normalized in [1,1][-1,1].
@@ 6.1 Quantitative & Qualitative Comparisons
Table 1: Quantitative comparisons of camera trajectories on the MultiCamVideo [2] dataset. Average results of all video sequences are reported. $\dagger$: To avoid memory overflow, we sample images with interval=8 for VGGT [37]; by default, other methods use all frames. The right side shows one reconstruction process of our method.
@@
Table 2: Quantitative comparisons of camera trajectories on the MultiCamRobolab dataset. Average results of all video sequences are reported. $\dagger$: To avoid memory overflow, we sample images with interval=8 for VGGT [37]; ‘X’ means fail to reconstruct scenes: not all images are successfully registered. GPU memory reports the peak inference memory consumption.
@@
First, we report the camera pose evaluation on the MultiCamVideo dataset and MultiCamRobolab in Tab.˜1, Tabs.˜2 and 3. Our method achieves the best results in the MultiCamVideo datasets, as shown in Tab.˜1. The right side in Tab.˜1 shows the visualization of one reconstruction process of our method. Our method achieves the overall best results in MultiCamRobolab datasets (shown in Tabs.˜2 and 3). Specifically, the traditional method COLMAP [33] can not handle such dynamic scenes well, which will generate noisy pose estimation. COLMAP can also fail in the non-overlap scenes (see the $RoboDog_{\text{non-overlap}}$ in Tab.˜2). VGGT cannot process all frames, even on a single A100-40GiB machine. For other feed-forward models, Fast3R also can not handle dynamic objects well; though CUT3R is designed for processing videos, it can not achieve satisfactory results (see Fig.˜4, for each separate video clip, CUT3R can not generate smooth results). In comparison, FastVGGT’s results are the best in all feed-forward models, which reveals the advantage of processing all frames together instead of test-time optimization used in CUT3R. In addition, the results indicate that FastVGGT is robust to dynamic objects in the scene. This observation motivates further investigation into the conditions under which such reconstruction models succeed or fail in dynamic environments. Our method achieves the best overall performance across all scenes (while consuming the least GPU memory), except for $RoboDog_{\text{non-overlap}}$. In this case, the cameras have no overlapping fields of view, causing the spatio-temporal connections to degenerate into purely temporal connections within each individual camera stream (more discussions can be found in the ablation study Sec.˜6.2).
@@
Table 4: Quantitative comparisons of frame depth and scene consistency on the MultiCamRobolab dataset. Average results of all video sequences are reported. $\dagger$:To avoid memory overflow, we sample images with interval=8 for VGGT [37]; by default, other methods use all frames.
@@
Input videos Ours FastVGGT [34] CUT3R [39] Fast3R [43]

### 14 Mar 2026 v2

@@ 5.1 Baseline methods
We select COLMAP [schonberger2016structure] as the baseline to represent the classical optimization-based structure-from-motion method. Since COLMAP only generates sparse point models, we limit our comparison to camera pose evaluation when using COLMAP. In experiments, we set the known camera intrinsics for COLMAP.
We also evaluate against recent state-of-the-art feed-forward models: Fast3R [yang2025fast3r], VGGT [wang2025vggt], and a memory-efficient version of VGGT FastVGGT [shen2025fastvggt]. These methods demonstrate strong performance in scenarios where traditional methods often fail. Finally, we include CUT3R [wang2025continuous] as a baseline due to its specific design for video processing, while it also supports unstructured image inputs. When running CUT3R, we sequentially feed frames from each camera video (i.e., camera-1 followed by camera-2) to fully leverage its temporal processing capabilities.
@@ 5.2 Metrics
We evaluate depth quality using Absolute Relative Depth (Abs Rel) and Delta accuracy [li2025megasam, zhang2024monst3r]($\delta<1.25$, i.e., the percentage of predicted depths within a 1.25-factor of true depth). Per-frame depth evaluation is conducted by scaling the predicted depth and offset to the ground truth.
@@ 5.3 Datasets
In addition to our self-collected MultiCamRobolab dataset, we also evaluate our method against the MultiCamVideo-Dataset [bai2025recammaster]. MultiCamVideo-Dataset is a multi-camera synchronized video dataset rendered in simulation using Unreal Engine, which includes synchronized multi-camera videos and their corresponding camera trajectories. Each video clip has 81 frames with 10 different camera perspectives. In our evaluation, we randomly extract 50 clips and 3 random cameras for each clip in the MultiCamVideo-Dataset. All images are cropped to $512\times 384$, and trajectories are normalized in [1,1][-1,1].
@@ 6.1 Quantitative & Qualitative Comparisons
Table 1: Quantitative comparisons of camera trajectories on the MultiCamVideo [bai2025recammaster] dataset. Average results of all video sequences are reported. $\dagger$: To avoid memory overflow, we sample images with interval=8 for VGGT [wang2025vggt]; by default, other methods use all frames. The right side shows one reconstruction process of our method.
@@
Table 2: Quantitative comparisons of camera trajectories on the MultiCamRobolab dataset. Average results of all video sequences are reported. $\dagger$: To avoid memory overflow, we sample images with interval=8 for VGGT [wang2025vggt]; ‘X’ means fail to reconstruct scenes: not all images are successfully registered. GPU memory reports the peak inference memory consumption.
@@
First, we report the camera pose evaluation on the MultiCamVideo dataset and MultiCamRobolab in Tab.˜1, Tabs.˜2 and 3. Our method achieves the best results in the MultiCamVideo datasets, as shown in Tab.˜1. The right side in Tab.˜1 shows the visualization of one reconstruction process of our method. Our method achieves the overall best results in MultiCamRobolab datasets (shown in Tabs.˜2 and 3). Specifically, the traditional method COLMAP [schonberger2016structure] can not handle such dynamic scenes well, which will generate noisy pose estimation. COLMAP can also fail in the non-overlap scenes (see the $RoboDog_{\text{non-overlap}}$ in Tab.˜2). VGGT cannot process all frames, even on a single A100-40GiB machine. For other feed-forward models, Fast3R also can not handle dynamic objects well; though CUT3R is designed for processing videos, it can not achieve satisfactory results (see Fig.˜4, for each separate video clip, CUT3R can not generate smooth results). In comparison, FastVGGT’s results are the best in all feed-forward models, which reveals the advantage of processing all frames together instead of test-time optimization used in CUT3R. In addition, the results indicate that FastVGGT is robust to dynamic objects in the scene. This observation motivates further investigation into the conditions under which such reconstruction models succeed or fail in dynamic environments. Our method achieves the best overall performance across all scenes (while consuming the least GPU memory), except for $RoboDog_{\text{non-overlap}}$. In this case, the cameras have no overlapping fields of view, causing the spatio-temporal connections to degenerate into purely temporal connections within each individual camera stream (more discussions can be found in the ablation study Sec.˜6.2).
@@
Table 4: Quantitative comparisons of frame depth and scene consistency on the MultiCamRobolab dataset. Average results of all video sequences are reported. $\dagger$:To avoid memory overflow, we sample images with interval=8 for VGGT [wang2025vggt]; by default, other methods use all frames.
@@
Input videos Ours FastVGGT [shen2025fastvggt] CUT3R [wang2025continuous] Fast3R [yang2025fast3r]